# Anomaly Induced Inflation in the Minimal Quartic Extension of Einstein’s Gravity

### v1

@@ 3.4.1 Perturbations around the de Sitter solution
Hence, the instability depends on the sign of the $\square R$ term in the trace of energy-momentum tensor. For the particle content of MSM, $\delta$ is negative, while MSSM gives positive values. As discussed above, the renormalization scheme dependence of this term makes it possible to adjust it to any value by introducing a counter $R^{2}$-term in Lagrangian. In most of the analysis, however, we simply consider the MSM case.
Inflation ends when the perturbation $H_{0}\left|\Delta H\right|$ is of the order of $H_{0}^{2}$, namely, with $A_{0}$ being a constant, $A_{0}\exp{\left(\xi H_{0}t_{f}\right)}\simeq H_{0}$, then the time when the inflation ends, $t_{f}$, can be estimated as
$t_{f}\simeq\frac{\zeta}{\xi H_{0}},$ (49)
where we have defined a new parameter $\zeta$ as
$\zeta=\log{\left(\frac{H_{0}}{\left|A_{0}\right|}\right)},$ (50)
since it will appear in the results frequently. If inflation is governed by a (quasi) de Sitter solution where $a(t)\simeq\exp{\left(H_{0}t\right)}$, we can introduce the number of e-folds $\mathcal{N}$ as
$\mathcal{N}\equiv\log{\left(\frac{a_{f}}{a_{i}}\right)}=H_{0}t_{f}.$ (51)
Moreover, we can define the effective EoS parameter for the effective fluid whose energy density and pressure is defined in (34).
$\omega_{\text{eff}}=\frac{p_{\text{eff}}}{\rho_{\text{eff}}}.$ (52)
For the de Sitter solution, $\omega_{\text{eff}}=-1$, but by introducing perturbations $H(t)=H_{0}+\Delta H(t)$, we obtain
$\omega_{\text{eff}}=-1-\frac{2\Delta\dot{H}}{3H^{2}}\simeq-1-\frac{2\xi A_{0}e^{\xi H_{0}t}}{3H_{0}}+\mathcal{O}(A_{0}^{2}).$ (53)
The accelerated expansion requires $w_{\text{eff}}<-1/3$. If the acceleration ends at $t_{f}$ in (49), by using the solution (45) and the number of $e$-folding $\mathcal{N}$, it can be said that during the inflation, the value of effective EoS changes between
$-1<\omega_{\text{eff}}\leq-1+\frac{2\zeta}{3\mathcal{N}}\,.$ (54)
Hence, to stop the acceleration at the end, i.e. $\omega_{\mathbf{eff}}(t_{f})\geq-1/3$, we found that the sign of $A_{0}$ is negative and the value should be $|A_{0}|\ll H_{0}$. Hence, the Universe starts from nearly de Sitter stage.
3.5 Observable constraints
Let us study some of the observables in the quartic theory. From the second time derivative of $a(t)$,
@@ 3.5 Observable constraints
To have viable inflation with a graceful exit, the slow-roll requires that the change in $\epsilon$ should also be small, so that we have another parameter $\eta$ to measure it.
@@
These two parameters, referred to as slow-roll parameters, play a pivotal role in analyzing the feasibility of inflation within the framework of standard inflationary models governed by scalar fields. In terms of these parameters, the current cosmological data constrain the spectral index $n_{s}$ and the tensor-to-scalar ratio $r$, which are given as
$n_{\mathrm{s}}=1-6\epsilon+2\eta\quad\text{and}\quad r=16\epsilon.$ (58)
Hence, in our model, we have the slow-roll parameters
$\epsilon\simeq-\frac{\Delta\dot{H}(t)}{H_{0}^{2}}=-\xi\frac{A_{0}}{H_{0}}e^{\xi H_{0}t}\,,$ (59)
$\eta\simeq\epsilon-\frac{\dot{\epsilon}}{2\epsilon H_{0}}=-\frac{\xi}{2}\left(1-\frac{A_{0}}{H_{0}}e^{\xi H_{0}t}\right)\,,$ (60)
and the spectral index
$n_{s}=1+\xi\left(7\frac{A_{0}}{H_{0}}e^{\xi H_{0}t}-1\right)+\mathcal{O}(A_{0}^{2})$ (61)
in terms of the model parameters. Since $|A_{0}|\ll H_{0}$, and $t\ll t_{f}$ during the inflation, we can express the spectral index approximately in a familiar form in terms of the number of $e$-folding,
$n_{s}\simeq 1-\frac{\zeta}{N}\,.$ (62)
Accordingly, the tensor-to-scalar ratio is given by
$r\simeq\frac{16\zeta e^{-\zeta}}{N}\,.$ (63)
To check if the model is consistent with the observations, we provide the observational constraints from the PLANCK satellite. According to the latest results [20] published in 2020,
$n_{s}=0.9649\pm 0.0042\;(68\%\text{CL})\quad\text{and}\quad r<0.101\;(95\%\text{CL})\,.$ (64)
The universe requires at least $N=60$ $e$-foldings during the inflation phase to address chronic problems such as flatness and horizon problems. Therefore, using 62 and according to Planck satellite results, we find that $\zeta\approx 2.4$. This yields the initial perturbation $A_{0}=H_{0}e^{-2.4}$, which is as small as required, and $\xi=0.04$ can be determined. Then, utilizing the second line in 47, we can establish the relationship between $\lambda$ and $\gamma$ within the allowed region of observations. In Figure 3, for instance, we depict the $\lambda$ vs $\gamma$ relationship and observe that for a small $\gamma<1$, there exists a value of $\lambda<1/4$ such that it falls within the physical range ($\kappa>0$) and is consistent with Planck results.
Figure 3: The relation between the effective dimensionless cosmological constant $\lambda$ vs the Born-Infeld parameter $\gamma$ that is consistent with Planck results ($n_{s}=0.96$) and in $c=\hbar=\kappa_{0}=1$ units.
In addition, with these parameter settings for example, the tensor-to-scalar ratio is found as $r\approx 0.058$ from the equation 63, which also satisfies observational constraints.

### v2

@@ 3.2.2 Perturbations around the de Sitter solution
Hence, the instability depends on the sign of the $\square R$ term in the trace of energy-momentum tensor. For the particle content of MSM, $\delta$ is negative, while MSSM gives positive values. As discussed above, the renormalization scheme dependence of this term makes it possible to adjust it to any value by introducing a counter $R^{2}$-term in Lagrangian since having $\Box R$ term in the trace of an energy-momentum tensor and having $R^{2}$ term in the Lagrangian are equivalent. In Appendix A, we investigate how having this affects the “potential term” of the equivalent point-like Lagrangian of our quartic $\mathcal{F}(R,\mathcal{G})$ theory (18). In most of the analysis, we simply consider the MSM case for the value of $\beta$, and examine negative values of $\delta$ for unstable initial de Sitter phase.
3.3 Dynamics of inflation
3.3.1 Dynamical system analysis
We first analyze our main differential equation (40) with a phase portrait analysis. This provides us general behavior of the solutions in the phase space without solving the equation. Using a change of variables,
$x=H(t)\quad\text{and}\quad y=\dot{H}(t),$ (46)
it can be written as a first order differential equation
With our definitions of $x$ and $y$, the velocity field reads
$(v_{x},v_{y})=\left(\frac{dx}{dt},\frac{dy}{dt}\right)=(\dot{H},\ddot{H})=(y,y\,f(x,y))\,.$
Then, plotting this velocity field, one can find and analyze the critical points and their characteristics of the system in $xy$-plane. In Fig.(1), we provided the phase portrait for the Minimum Standard Model (MSM) in four different regimes.
(a) Einstein’s gravity with trace-anomaly. ($\gamma=0$)
(b) Minimal quartic model without trace-anomaly. ($\gamma=0.0025$)
(c) Minimal quartic model with trace-anomaly. ($\gamma=0.0025$)
(d) Minimal quartic model with trace-anomaly. ($\gamma=0.0060$)
Figure 1: The phase-portrait of the system described by the EoM (40) for the MSM case with parameter settings $x=H(t)$ and $y=\dot{H}(t)$. The figure is plotted in $c=\hbar=\kappa_{0}=1$ units and the value of the bare cosmological constant is chosen arbitrarily as $\Lambda_{0}=6$.
In the first scenario depicted in Fig.(1-a), we examine the case where there is no contribution from the modified gravity, representing a pure trace-anomaly situation within the framework of Einstein’s gravity. Two critical points can be seen along the $y=0$ axis, corresponding to two distinct de Sitter solutions. For the particle content of MSM, the second solution is identified as unstable. Consequently, the universe is presumed to have originated nearly from this point, having a negative $\dot{H}$, subsequently transitioning towards the first critical point, which acts as the final attractor of the system, determined by $\Lambda_{0}$, and may correspond to the late-time acceleration. In the absence of an intrinsic cosmological constant, the ultimate attractor point resides at $(x,y)=(0,0)$. Thus, during the inflationary epoch, the Hubble parameter $H$ experiences a slow decrease while maintaining $\dot{H}$ in near constancy, eventually oscillating around the final attractor, indicative of the reheating era.
In the second scenario, depicted in Fig.(1-b), we switch off the trace anomaly and investigate the dynamics of pure modified gravity. Once again, two distinct de Sitter solutions are evident along the $y=0$ axis. Additionally, a solid black line is introduced along the x-axis to indicate a non-physical region where $\lambda>1/4$, a characteristic feature typical of Born-Infeld-type theories, setting an upper limit on the effective cosmological constant. Consequently, the second solution is nonphysical, as it intersects with the solid line, resulting in the existence of a unique de Sitter solution. In this context, a critical surface manifests along the $y=0$ axis, changing the characteristics of critical points both below and above this surface. The unique de Sitter solution remains stable within the domain where $\dot{H}<0$, yet transitions to instability when $\dot{H}>0$ is encountered. Furthermore, the critical surface is identified as unstable within these points, while exhibiting attractor behavior beyond this region.
In Fig.(1-c), we explore our modified gravity theory gently introducing a small value for $\gamma$ in the presence of the trace anomaly. In comparison to the first scenario, the introduction of modified gravity results in the emergence of three critical points situated above the de Sitter solutions observed in the original trace-anomaly case. However, one of these critical points exhibits a notably high $\dot{H}$ value, surpassing the upper limit of this plot and consequently playing no role in the inflationary solutions. The remaining two critical points reside above the critical points along the $y=0$ axis and are positioned on a critical surface represented by a curve that converges toward the x-axis as $H\rightarrow\infty$. Notably, the critical surface with critical points on it in pure modified gravity case is raised within the $\dot{H}>0$ region, while still maintaining the de Sitter solution due to the trace anomaly within the physically plausible range, $\kappa>0$. Under an initial condition where $\dot{H}\lesssim 0$, the dynamics of inflation remain largely unaffected by the modified gravity.
In Fig.(1-d), to understand the effect of modified gravity on trace-anomaly-driven inflation, we play with the value of $\gamma$. Increasing $\gamma$ makes the top critical points get closer to the de Sitter points. The solid line indicates the $\kappa<0$ region and, as $\gamma$ increases, it includes the initial de Sitter point which means inflation cannot happen in this (unphysical) case. This imposes an upper bound on $\gamma$ so that the effects of modified gravity should not exceed the trace anomaly for a viable inflationary model.
(a) Trace-anomaly in Einstein’s gravity with MSSM particle content
Figure 2: The phase-portrait of the system described by the EoM (40) for MSSM case. We shade off the region $|y/x^{2}|>1$ indicating the outside of the slow-roll regime which is not preferred for initial conditions. The curves highlighted in color in (b), are some example solutions with (green) and without (red) graceful exit. The figure is plotted in $c=\hbar=\kappa_{0}=1$ units and the parameters are set as $\gamma=0.004$ and $\Lambda_{0}=6$.
Furthermore, we investigate the scenario wherein the sign of the $\Box R$-term is positive, indicative of an initial stable point. This condition finds representation in the context of MSSM as an illustrative example model encompassing the particle content pertinent to this case. In Fig.(2), the phase-portrait of the system is plotted without and with the effects of modified gravity for the MSSM case respectively, and from Fig.(2-a), it can be seen that the second de Sitter solution due to the trace-anomaly is a stable critical point. In this case, a graceful exit from Inflation requires its mechanism to disrupt this state. In Ref. [20], a possible transition from stable to unstable phases of inflation is discussed, and different quantum mechanisms that may produce large $R^{2}$-term during the inflation are proposed for a graceful exit, for example. On the contrary, we show that the minimal extension of Einstein’s gravity makes possible an exit from inflation without any other mechanism. In Fig.(2-b), we have plotted the contribution of modified gravity in this case. As in MSM case, modified gravity creates additional critical points on the phase-portrait, visible ones of those in the plotted region are labeled as $P_{i}(x_{i},y_{i})$ in the graph, but this time the two of them lying on a critical surface are in $\dot{H}<0$ region. Note that increasing the intensity of modified gravity makes this critical surface closer to the $y=0$ line. Therefore, if the universe starts with $H\simeq x_{1}\simeq x_{2}$ and a small value of $\dot{H}$ from below this surface, $\dot{H}\lesssim y_{2}$ for example, then the exit from inflation can occur due to the attractor behavior of the surface between 00 and $x_{3}$. This prevents $\dot{H}$ from flowing through negative infinity and ends inflation with $H\lesssim x_{3}$. If the strength of modified gravity is high enough, the initial point for $\dot{H}$ can be in the slow-roll regime, i.e. $\left|\dot{H}/H^{2}\right|\ll 1$ while keeping $\kappa$ positive. In Fig.(2-b), the green curves are some examples of possible inflationary solutions with graceful exit while the red ones does not have enough initial $\dot{H}$ value so that the graceful exit is not possible for those similar to the Einstein’s gravity. We should note that the necessary number of e-folding is not guaranteed for all of these solutions, but a particular realistic solution with enough number of e-folding is analyzed in the next section.
3.3.2 Numerical solutions
Here, we numerically solve EoM which is the $00-$component of the field equations given explicitly in (40) for different numerical values of $\gamma$ and $\delta$. Recall that the value of the parameter $\beta$ is fixed by renormalization, but we have a freedom to set $\delta$ to any desired value by adding $R^{2}$ counter term to the action due to its renormalization scheme dependence. Moreover, because the parameter $\Lambda_{0}$ is related to the final attractor of the system, we set it to zero $\Lambda_{0}=0$ for the rest of the analysis, but it can be set to a value for unified early and late time acceleration model. Hence, our main free parameters are BI parameter $\gamma$ and the coefficient of $\Box R$ term, $\delta$, to fit the solution into desired shape. In Fig.(3), we give four numerical solution of the field equation for $H(t)$ to judge the model parameters $\gamma$, $\Lambda_{0}$, and $\delta$, with the parameter $\beta$ set to $\beta=253/(1440\pi^{2})$ for the content of the MSM particles.
Figure 3: Numerical solutions of $00$-component of field equations for $H(t)$ with different parameter settings in $c=\hbar=\kappa_{0}=1$ units. Here, the value of $\beta=253/(1440\pi^{2}$ is given by MSM particle content. The other parameters that we used for numerical solutions are given as
$1$ $\gamma=0.0043$, $\delta=-0.25$, $\Lambda_{0}=0$; $2$ $\gamma=0.0043$, $\delta=-0.35$, $\Lambda_{0}=0$; $3$ $\gamma=0.0025$, $\delta=-0.25$, $\Lambda_{0}=0$; $4$ $\gamma=0.0043$, $\delta=-0.25$; $\Lambda_{0}=0.2$. The initial conditions for the numerical solutions are taken as $H(0)=H_{i}-10^{-5}$ and $\dot{H}(0)=0$ where $H_{i}$ is the solution of (41) which is the value of initial de Sitter phase.
With the help of the numerical results, we have a chance to examine the inflation scenario that we revealed in the phase portrait analysis in more detail. Then, the typical solution of $H(t)$ can be described as follows. $H$ starts with an initial value of the unstable vacuum state determined by $\beta$ and $\gamma$ as a solution of (41), and slowly roll down from this value until it enters a linearly decreasing phase, $H\propto\ -t$, through to zero which is a typical behavior of $R^{2}$ gravity in the slow roll regime [21]. Then, bouncing back from zero, $H$ approaches the value of the final vacuum determined by $\Lambda_{0}$ by oscillations with decreasing amplitude called the reheating era. If $\Lambda_{0}=0$, the final value of $H\rightarrow 0$. In Fig.(4), one of the solutions can be seen in the phase space that we defined in the previous section. In addition, using the numerical solution of $H(t)$, we have also plotted the function $a(t)$ in Fig.(5), from which the behavior of the scale factor during inflation can be seen as quasi-exponential while during the reheating phase it can be approximated by $a(t)\propto t^{\frac{2}{3}}$.
Figure 4: The plot of the numerical solution $1$ in $H-\dot{H}$ phase space. The solution is consistent with the case of phase portrait (1)(c) that we discussed in the dynamical system analysis.
(a) during inflation
(b) during reheating
Figure 5: Evolution of the scale factor $a(t)$ in time. In left panel (a), we see a quasi-exponential expansion which expands space about $10^{59}$ order. On the right panel (b), the behavior of $a(t)$ during reheating era is presented and shows that the scale factor evolves like matter dominated phase as $a\sim t^{2/3}$. Here, $t_{r}$ is the initial time of reheating and we re-scaled the initial value by normalizing with $a(tr)$ for this plot.
In our numerical analysis, the aim is to examined the role of parameters $\delta$ and $\gamma$ on the dynamics of inflation and the number of $e$-folding $\mathcal{N}$. At the end of inflation $e$-folding number $\mathcal{N}$ is given by
$\mathcal{N}\equiv\log{\left[\frac{a(t_{f})}{a(t_{i})}\right]}\,,$ (49)
where $t_{i}=0$ and $a(t_{i})=1$ at the initial time, and $t_{f}$ is the end of inflation where the Universe enters the oscillatory phase. By varying $\delta$ and $\gamma$ in many numerical solutions, in Fig.(6), we find the relations between $\mathcal{N}$ at the end of inflation and model parameters. Thus, the following observations can be made. First, the duration of the inflation is determined mainly by the term $\delta$, but $\gamma$ also contributes little. A higher $\gamma$ reduces the initial value of the $H$ since it enters the solution of de Sitter equation (41) and also it suppress the duration of inflation. However, for fixed $\beta$, there is a maximum value of $\gamma$ at which $\lambda=1/4$. For MSM particle content, $\beta=253/(1440\pi^{2})$, the highest value is $\gamma\simeq 0.00431$ so that a $\gamma$ with higher than that value makes $\kappa$ negative, which is undesired.
Figure 6: The dependence of total e-folding number $\mathcal{N}$ at the end of inflation on the parameters $\delta$ and $\gamma$. The data points are found by varying model parameters and performing numerous numerical solutions. In (a), the relation with respect to $\delta$ is calculated at fixed $\gamma=0.0043$, while in (b), the relation is at fixed $\delta=-0.25$.
The slow-roll parameters
We can also check if slow-roll parameters during the inflation are small enough. From the second time derivative of $a(t)$,
@@ 3.3.2 Numerical solutions
To have a viable inflation with a graceful exit, the slow-roll requires that the change in $\epsilon$ should also be small, so that we have another parameter $\eta$ to measure it.
@@
These two parameters, referred to as the Hubble slow-roll parameters, play a pivotal role in analyzing the feasibility of inflation within the framework of standard inflationary models. In Einstein’s gravity with a single scalar field, we can express some of the observables such as spectral index of scalar curvature perturbations and tensor-to-scalar ratio in terms of these parameters. In the modified gravity case, however, there could be more than two slow-roll parameters depending on the structure of the gravity model. In that case, one can introduce parameters as needed, such as [22],
$\epsilon_{1}=-\frac{\dot{H}}{H^{2}},\quad\epsilon_{2}=\frac{\ddot{\mathcal{G}}}{H\dot{\mathcal{G}}},\quad\epsilon_{3}=\frac{\dot{f}_{R}}{2Hf_{R}},\quad\epsilon_{4}=\frac{\dot{E}}{2HE}\quad\text{with}\quad E=\frac{3\dot{f}_{R}^{2}}{2\dot{\mathcal{G}}^{2}}\,$ (53)
and extend this set further if necessary. Then, the true relations between the slow-roll parameters and the observables come from the cosmological perturbations in that theory with scalar vector tensor (SVT) decomposition.
Here, we should make an important remark. Our theory, just like the model of Starobinsky, has a massless spin-2 and a massive spin-0 field, the latter coming from the quantum anomaly. (For the particle content of the model, see Appendix B and also [24].) The mass of the scalaron in the Starobinsky case ($\mathcal{L}=R+\kappa_{0}\alpha R^{2}$), is $m^{2}_{\phi}=\frac{1}{6\kappa_{0}\alpha}$, while our quartic extension contributes to it via effective gravitational constant so that in the maximally symmetric background, it becomes $m^{2}_{\phi}=\frac{1}{6\kappa\alpha}$ where $\kappa$ is given by (5).
In Fig. (7-b), we have seen that all of these parameters are small during the inflation so that inflationary solutions within this theory fulfill the slow-roll conditions. If the horizon crossing happens at $\mathcal{N}_{*}=50-60$, then the order of these parameters as power of $10$ are given as $\mathcal{O}\left(\eta,\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}\right)=\{-1,-3,-1,-1,-3\}$. Hence, in the linear approximation with suitable chosen model parameters, any linear combination of these parameters in the expression of spectral index with the coefficients $c_{i}$, coming from perturbation study, such as $n_{s}=1-\sum_{i=1}^{n}c_{i}\epsilon_{i}$ can satisfy observational constraints [23] if $c_{i}$’s are smaller or around order of 10.
(a) Number of $e$-folding
(b) slow roll parameters during inflation
Figure 7: The number of $e$-folding during the inflation with respect to cosmic time t is given in (a) for different solutions presented in Fig. (3). In (b), The evolution of slow roll parameters $\epsilon_{i}$ and $\eta$ during the inflation is given with respect to number of $e$-folding $\mathcal{N}$ for the solution $1$.
Approximate solution
The numerical solutions show that the typical behavior of the function $H(t)$ during inflation is a sigmoid curve, which can be approximated by using a generalized logistic function $\bar{H}(t)$ of the form
$\bar{H}(t)=\frac{H_{m}}{\left(1+e^{k(t-t_{0})}\right)^{n}},$ (54)
where $H_{m}$, $k$, $t_{0}$ are constants needs to be determined and $n$ is integer. For the numerical solution presented in Fig.(3) with model parameters
[…]
(a) $H$ vs $t$
[…]
3.4 Evolution of matter perturbations
The theory deserves the study of cosmological perturbations in full detail; however, since this is rather cumbersome work, let us make a numerical analysis following the work [25]. Consider the equation which governs the evolution of the matter fluctuations in the linear regime
$\ddot{\delta}_{m}+2H\dot{\delta}_{m}-\frac{1}{2}\kappa_{\mathrm{eff}}\rho_{m}\delta_{m}=0$ (55)
where $\rho_{m}$ is the matter density and $\kappa_{\text{eff }}$ is the effective gravitational coupling constant which, in $\mathcal{F}(R,\mathcal{G})$ gravity case, is
$\kappa_{\mathrm{eff}}=\frac{\kappa_{0}}{\mathcal{F}_{R}(R,\mathcal{G})}$ (56)
where $\kappa_{0}$ is the Newton gravitational constant. Notice that for maximally symmetric background, $\kappa_{\mathrm{eff}}=\kappa$ in (5). Here, we are considering perfect fluid matter that is added in action (19) as minimally coupled. We use EoM (34) with matter density contribution as
$\rho_{m}=\frac{3H^{2}}{\kappa_{0}}-\rho_{\mathrm{eff}}$ (57)
where $\rho_{\mathrm{eff}}$ includes contributions of both modified gravity, $\rho_{\mathrm{MG}}$, and trace-anomaly, $\rho$ in (38).
Substituting these into (55), and by using approximate solution of $H=\bar{H}(t)$ as a background expansion, one can find a numerical solution for the matter fluctuations $\delta_{m}(t)$. With the parameters chosen in previous section, the evolution of the matter perturbations during the inflation is given in Fig.(9)
(a) matter perturbations with respect to scale factor
Figure 9: Evolution of the matter perturbations during the inflation in with respect to scale factor. The matter perturbations are linearly grows during the initial phase, and almost frozen at the last stage. The Log-Log plot is in base $e$.
We observe linear growth in matter perturbations concerning the scale factor during the initial phase. Perturbations reach a maximum value, about a factor of $10^{9}$ in our case, and almost freeze by the end of inflation. This gives an overview of the matter perturbations, but a full cosmological perturbation theory with standard SVT decomposition in the metric should be studied for more accurate results.